OpenStudy (anonymous):

Write the equation of the line, in slope-intercept form, that contains the points (-1, -2) and (1, 2).

OpenStudy (misty1212):

what did you get for the slope?

OpenStudy (anonymous):

2

OpenStudy (misty1212):

yes !

OpenStudy (misty1212):

so now what?

OpenStudy (anonymous):

That's why I'm asking..lol

OpenStudy (misty1212):

do you know the point slope formula?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

@Bookworm14

OpenStudy (bookworm14):

ok so were just gonna do the same thing we did a minute go: y-y1=m(x-x1)............... (y1 is your y intercept, m is your slope, x1 is your x intercept) so you have a slope of 2, your first y intercept is -2, and your first x intercept is -1, now just substitute those into the point-slope equation y- (-2) = 2(x- (-1) now rewrite it --> y+2= 2(x+1) now distribute --> y+2= 2*x + 2*1 now rewrite again --> y+2= 2x +2 subtract "2" from the "y" side, what you do on one side, do to the other -2 -2 y= 2x (this is now in slope- intercept form or y=mx+b)

OpenStudy (bookworm14):

@xo_kansasprincess_xo